ZoeyBiocth16:

a six sided die is rolled. what is the probability that the roll is 1 or 2

BlazeNeedsHelp:

combined it would be and 2 out of six chance

Pixel:

if it is rolled once, it has 6 sides, can land on 2 = 2/6 = 1/3

ZoeyBiocth16:

wow thank you

BlazeNeedsHelp:

yes or if you simplify a 1/3 chance

ZoeyBiocth16:

another way i did it was adding 1/6 +1/6 just wanted to make sure i was right

BlazeNeedsHelp:

yes you were good job

ZoeyBiocth16:

lmao thanks sir